Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

letters and envelopes - probability

Expert replies
by DavoodBeater » Mon Jan 19, 2009 8:27 am
Please explain. Thanks.

Tanya prepared 4 different letters to be sent to 4 different addresses. For each letter, she prepared an envelope with its correct address. If the 4 letters are to be put into the 4 envelopes at random, what is the probability that only 1 letter will be put into the envelope with its correct address?

OA: [spoiler]1/3[/spoiler]
Join the discussion
Source: — Problem Solving |

Re: letters and envelopes - probability

by logitech » Mon Jan 19, 2009 9:26 am
A B C D
1 2 3 4

Let's say only A1 is matched correctly.

For 2 can choose only C or D
For 3 can choose only D or B
For 4 can choose only B or C

So actually for every letter that is in its correct address(A1) you can have two different ways to distribute the other three to ALL incorrect addresses.

Since we have 4 letters

4x2= 8 Desired Scenarios

4 letter can be sorted as 4! ways

8/4! = 1/3
LGTCH
---------------------
"DON'T LET ANYONE STEAL YOUR DREAM!"
Join the discussion

by sachinkr » Mon Jan 19, 2009 9:28 am
Let's Assume 4 letters are labeled a,b,c,d and envelope are labeled A,B,C,D.

Total number of ways 4 letters can be put in 4 envelopes = 4*3*2*1 = 24

Number of ways only 1 letter is in correct envelope

In case label 'a' letter is put in label A envelope. the number of cases are :
1: A (a), B (c) , C (d), D (b)
2: A (a), B (d) , C (b), D (c)

Similarly b,c,d letters can be placed in their respective envelope. So in all there are 4 * 2 = 8.

So the probability is 8/24 => 1/3.
Join the discussion

by DavoodBeater » Mon Jan 19, 2009 10:12 am
Thanks a lot for your answers.
Join the discussion

by amitabhprasad » Mon Jan 19, 2009 10:20 am
Four letter say labeled as A,B,C,D.
P(A) is in the right envelop = 1/4
P(B) is in the right envelop = 1/3
==> P(B) is not in the right envelop = 2/3
P(C) is int he right envelop = 1/2
==> P(C) is not in the right envelop = 1-1/2 = 1/2
==>P(D) not in the right envelop = 1
Thus probability of only getting "A" in the right envelop =
P(A) && P(!B) && P(!C) && P(!D) = P(A)*P(!B)*P(!C)*P(!D) = 1/4*2/3*1/2*1 = 1/12
Similarly probability "B"of P(B) = 1/12 ; P(C) = 1/12;P(D) = 1/12

We have 4 envelop chance of any one of them is wrong =
P(A or B or C or D) = P(A)+P(B)+P(C)+P(D) = 1/12+1/12+1/12+1/12
=4/12 = 1/3
Join the discussion

by DavoodBeater » Mon Jan 19, 2009 11:48 am
I did the same.
but you know, the first success is not always the first one. So if you faced a failure at first, the chance of success would be 1/3. so;
3/4 * 1/3 * 1/2 * 1
but this way will not lead us to the same answer and I dont know why.
the whole answer which I did:

1/4 * 2/3 * 1/2 * 1 : first success
3/4 * 1/3 * 1/2 * 1 : second success
3/4 * 2/3 * 1/2 * 1 : third success
3/4 * 2/3 * 1/2 * 1 : last success
the sum is 17/24 not 8/24
Join the discussion